OpenStudy (anonymous):

Why is radical25 not equal to -5?

OpenStudy (lochana):

because 25 is not a variable.

OpenStudy (solomonzelman):

\(\normalsize\color{black}{ \sqrt{25} }\) can be equal to \(\normalsize\color{black}{ 5 }\) or \(\normalsize\color{black}{ -5 }\).

OpenStudy (lochana):

no , \[\sqrt{25}\] only can be equal for +5

OpenStudy (solomonzelman):

No

OpenStudy (solomonzelman):

\(\normalsize\color{black}{ \sqrt{25}=±5 }\) (Since, you know that \(\normalsize\color{black}{ (-5)^2=25 }\) and \(\normalsize\color{black}{ 5^2=25 }\) .

OpenStudy (solomonzelman):

You are confusing 2 different concepts. A square root of a number (if the number is positive) has 2 real values, like \(\normalsize\color{black}{ \sqrt{36}=±6 \color{white}{\Large()ccc}}\) \(\normalsize\color{black}{ \sqrt{25}=±5 \color{white}{\Large()ccc} }\) \(\normalsize\color{black}{ \sqrt{16}=±4 \color{white}{\Large()ccc}}\) \(\normalsize\color{black}{ \sqrt{9}~~~=±3\color{white}{\Large()ccc} }\) \(\normalsize\color{black}{ \sqrt{4}~~~=±2\color{white}{\Large()ccc} }\) \(\normalsize\color{black}{ \sqrt{1}~~~=±1\color{white}{\Large()ccc} }\)

OpenStudy (solomonzelman):

And the other concept that there is no such a thing as \(\normalsize\color{black}{ \sqrt{-b} }\)

OpenStudy (solomonzelman):

(in terms of a real number solution)

OpenStudy (lochana):

well may be you right, but my maths teacher taught me something like this. if x^2 = 25 then x =- or + \[\sqrt{25}\] and then x = - or + 5 this means \[\sqrt{25}\] = 5 thats why i'm disagree with you